# FisherRF: Active View Selection and Uncertainty Quantification for Radiance Fields using Fisher Information

### v1

@@ Abstract
This study addresses the challenging problem of active view selection and uncertainty quantification within the domain of Radiance Fields. Neural Radiance Fields (NeRF) have greatly advanced image rendering and reconstruction, but the limited availability of 2D images poses uncertainties stemming from occlusions, depth ambiguities, and imaging errors. Efficiently selecting informative views becomes crucial, and quantifying NeRF model uncertainty presents intricate challenges. Existing approaches either depend on model architecture or are based on assumptions regarding density distributions that are not generally applicable. By leveraging Fisher Information, we efficiently quantify observed information within Radiance Fields without ground truth data. This can be used for the next best view selection and pixel-wise uncertainty quantification. Our method overcomes existing limitations on model architecture and effectiveness, achieving state-of-the-art results in both view selection and uncertainty quantification, demonstrating its potential to advance the field of Radiance Fields. Our method with the 3D Gaussian Splatting backend could perform view selections at 70 fps. Source code and other materials are available at https://jiangwenpl.github.io/FisherRF/.
@@ 1 Introduction
Figure 1: A brief overview of our method Given a radiance field that was trained with a limited number of views, our method could find the next best view that could maximize information gain by computing the Fisher Information of the radiance field. We illustrate the information gain as a heat map on the viewing sphere and show four of the candidate views. Our model can quantify pixel-wise uncertainty, visualized on the bottom, by examining the Fisher Information on the related parameters of the ray.
@@
Neural Radiance Fields brought back image rendering and reconstruction from multiple views to the center of attention in the field of computer vision. Novel volumetric representations of radiance fields and differentiable volumetric rendering enabled unprecedented advances in image-based rendering of complex scenes both in terms of perceptual quality and speed. Neural Radiance Fields (NeRF) have recently garnered significant attention in the field of Computer Vision and have seen enhancements for faster rendering and training, larger-scale reconstruction, dynamic modeling, and generative capabilities. Recently, 3D Gaussian Splatting has demonstrated distinct advantages in real-time rendering and explicit point-based parameterizations without neural representations. However, the majority of approaches that estimate a radiance field need tens of viewpoints. When we are limited to a low number of 2D images, uncertainty naturally arises from occlusions, depth ambiguities, and aleatoric errors in the imaging process. It is crucial to establish a criterion for selecting information-maximizing views before knowing what the actual error will be in these views. Quantifying the observed information of a NeRF model is challenging, given that NeRF models are typically regression-based and scene-specific. The challenge intensifies when we aim to leverage quantified uncertainties for active view selection, especially when the selection candidates are only $SE(3)$ camera poses for acquiring new observations, a.k.a capturing new images.
@@
Previous approaches for uncertainty quantification and view selection can be broadly categorized into two groups: variational white-box models and predictive black-box models. White-box models integrate conventional NeRF architectures with Bayesian models, such as reparameterization [25, 21, 26] and Normalizing Flows [32]. Black-box methods, on the other hand, do not modify the existing model architecture but seek to quantify predictive uncertainty by examining the distribution of predicted outcomes [40, 5].
White-box models depend on specific model architectures and are often characterized by slower training times due to the challenges associated with probabilistic learning. Conversely, existing black-box models either focus solely on studying the distribution of densities along a ray, assuming that the density distribution on a ray should be as sharp as possible, or rely on Monte Carlo-style sampling techniques [38] to quantify the perturbations of their model. In this study, our primary objective is to quantify the observed information of a Radiance Field model and utilize it to select the optimal view with the highest information gain. To achieve this, we propose the use of Fisher Information, which represents the expectation of observation information. This quantity is directly linked to the second-order derivatives or Hessian matrix of the loss function involved in optimizing Radiance Field models.
[…]
@@
We implemented the computation of Fisher Information on top of two types of Radiance Field models: point-based 3D Gaussian Splatting [13] and Plenoxels [29]. In 3D Gaussians, we compute the Hessians of the mean, covariances, opacity, and spherical harmonics with respect to the negative log-likelihood. In Plenoxels, we compute the Hessians of density and spherical harmonics on the sparse voxel grid representation of Plenoxel with respect to its negative log-likelihood. When selecting candidates, we apply a greedy policy, choosing the view with the highest information gain. To make the computation of information gain tractable, we use an approximation of the decrease in entropy by its upper bound, which is the trace of the product of the Hessian of the candidate view times the Hessian with respect to all previous views. These matrices are highly sparse due to the disentangling of scene parameters with respect to disparate rays. The sparsity allows us a computation of the matrix trace above that is as cost-effective as back-propagation, enabling us to evaluate views at 70 fps when we use 3D Gaussian Splatting. We carried out an extensive evaluation in both active view selection and uncertainty quantification. Our evaluation contains the PSNR in the hold-out test as a function of the number of views and the selection regime (one or more views at a time) as well as an AUSE evaluation of the uncertainty and an evaluation of the stratification error. The quantitative and qualitative results unequivocally demonstrate that our approach surpasses previous methods and heuristic baseline by a significant margin. To summarize, our main contributions are as follows:
  • A novel quantification of uncertainty that we use in computing the information gain of candidate views.
@@
  • An efficient computation of the information gain exploiting the sparse structure of the scene rendering problem.
@@
  • We were able to compute a pixel-wise uncertainty quantification and visualization.
@@
  • Extensive experimental evaluation showing that our active view selection outperforms existing active approaches.
@@ 2 Related Works
Active Learning and Radiance Fields are prosperous research fields with various research directions. In this section, we limit our literature overview on works on view selection and uncertainty quantification for radiance fields. We refer the readers to literature reviews if they are interested in Radiance Fields [8, 7] or Deep Active Learning [28, 41]. Besides, Fisher Information has been extensively studied in deep active learning [2, 1, 17, 16]. Notably, Kirsch et al. [14] unified existing works in active learning for deep learning problems from the perspective of Fisher Information, which shares many insights with our method.
@@
In the Neural Radiance Fields, Chen et al. [5], Yan et al. [37] and Zhan et al. [40] attempted to quantify the uncertainty in a scene by the distribution of densities on a casted ray. Pan et al. [25] and Shen et al. [32, 31] designed Bayesian models by re-parametrizing the NeRF model. However, they only tackled the predictive uncertainty of the model that did not relate to the observed information of the parameters. Sunderhauf et al. [35] propose an additional uncertainty measure in uncharted regions, determined by the ray termination probability on the learned geometry. Regarding concurrent work, Goli et al. [9] also introduces Fisher information to quantify the uncertainty. However, they applied a considerable simplification by computing the Hessian over a hypothetical perturbation field, while we directly estimate the Fisher Information on the model parameters. Besides, the focus of their method is cleaning up noisy 3D positions in a NeRF model. Yan et al. [38] discussed the intuition of flat minimum and quantified the predictive uncertainty of neural mapping models through the lens of loss landscape. However, they only approximate the uncertainty of the variances from the output of neighboring points. Our method directly quantifies the observed information of the parameters by computing the Hessian matrix of the log-likelihood function of the model’s objective. To the best of our knowledge, we are the first method that quantifies the uncertainty of explicit radiance field models, thanks to the powerful framework of 3D Gaussian and our efficient formulation.
Active Learning for Radiance Field
Although the next best views selection problem was extensively studied before radiance fields gained popularity, there has not been much literature on active “training” view selection for novel view rendering. ActiveNeRF [25] studied the next best view selection and it is the closest approach to the view selection problem we are working on. Although Ran et al. [26], Yan et al. [38], Zhan [40] included the next best view selection in their neural mapping or reconstruction system, their goal was not to quantify the observed information of a radiance field.
@@ 3 Technical Approach
In this section, we first introduce the background knowledge of volumetric rendering in Sec. 3.1, which radiance field models, including our backbone model 3D Gaussian Splatting and Plenoxel, are widely used. We then introduce how we use Fisher Information in Sec. 3.2. By leveraging Fisher Information, we demonstrate how to select the next best view and next best batch of views in Sec. 3.3 and Sec. 3.4. Finally, we showcase that our method could quantify the pixel-wise uncertainties in Sec. 3.5.
3.1 Preliminaries: Volumetric Rendering and Gaussian Splatting
3D Gaussian Splatting [13], Plenoxels [29], and Neural Radiance Fields [22] all use volumetric rendering for image formation. Consider a camera ray $\text{r}(t)=\text{o}+t\text{d}$ originating from the camera center $\text{o}\in\mathcal{R}^{3}$ and passing through a specific pixel on the image plane. The color of the pixel can be expressed as:
[…]
$\hat{C}(\text{r})=\sum_{i=1}^{N_{s}}T_{i}\left(1-\exp(-\sigma_{i}\delta_{i})\right){\bf c}_{i}$ (2)
$\alpha_{i}=\text{exp}\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right)(1-\text{exp}(-\sigma_{i}\delta_{i}))$ (3)
Here, $\delta_{i}=t_{i+1}-t_{i}$ represents the distance between adjacent samples, and $N_{s}$ indicates the number of samples. Based on this approach, NeRF approaches optimize the continuous function $F_{\theta}$ by minimizing the squared reconstruction errors between the ground truth obtained from RGB images ${\mathcal{I}_{i=1}^{N}}$ and the rendered pixel colors. Although 3D Gaussian Splatting has a different rendering pipeline, the image formation is still similar, whereas, in 3D Gaussian Splatting, ${\bf c}_{i}$ is the color of each 3D Gaussian given the view direction ${\bf d}$ and $\sigma_{i}$ is given by evaluating a 2D Gaussian with covariance $\Sigma$.
@@ 3.2 Fisher Information in Volumetric Rendering
Fisher Information is a measurement of information that an observation $({\bf x},{\bf y})$ carries about the unknown parameters $w$ that model $p({\bf y}|{\bf x};{\bf w})$. In the problem of novel view synthesis, $({\bf x},{\bf y})$ are the camera pose ${\bf x}$ and image observation ${\bf y}$ at pose ${\bf x}$, respectively, whereas ${\bf w}$ are the volumetric parameters of the radiance field. The objective of neural rendering is equivalent to minimizing the negative log likelihood (NLL) between rendered images and ground truth images on the holdout set, which inherently represents the quality of scene reconstruction
@@
$\mathcal{I}({\bf w})=-\mathbb{E}_{{\bf y}|{\bf x}\sim p_{\bf w}}\big{[}\frac{\partial^{2}\log p({\bf y}|{\bf x},{\bf w})}{\partial{\bf w}^{2}}\big{|}{\bf w}\big{]}={\mathbf{H}}^{\prime\prime}[{\bf y}|{\bf x},{\bf w}]$ (5)
@@ 3.3 Next Best View Selection Using Fisher Information
$\mathcal{I}[{\bf w}^{*};\{{\bf y}^{acq}_{i}\}|\{{\bf x}^{acq}_{i}\},D^{train}]$
$=H[{\bf w}^{*}|D^{train}]-H[{\bf w}^{*}|\{{\bf y}^{acq}_{i}\},\{{\bf x}^{acq}_{i}\},D^{train}]$ (6)
@@
When the log-likelihood has the form of Eq. (4), in our case the rendering error, the difference of the entropies in the R.H.S. of Eq. (6) can be approximated as [14]:
@@
Figure 2: An example of the Jacobinas of the rendering equation on a casted ray. We plot the absolute values of the Jacobian of three channels in one ray w.r.t the sampled points on the ray in Plenoxel. Most sampled points in volumetric rendering only have very small values in their corresponding Jacobian matrix, prompting the sparsity of the Hessian matrix in Radiance Fields.
@@
This approximation also allows us to invert the Hessian matrix ${\mathbf{H}}^{\prime\prime}[{\bf y}|{\bf x},{\bf w}^{*}]$ easily. We further showcase the sparsity of the Jacobian and Hessian matrix in Fig. 2 and Fig. 3.
Figure 3: An example of the Hessian matrix on the parameters of Plenoxel. We compute the Hessian of the NLL function of volumetric rendering following the Eq. (10).
@@ 3.4 Batch Active View Selection
Selecting multiple views to capture new images is useful for its possible applications, such as view planning and scene reconstruction. If we directly use Eq. (8) to select a batch of acquisition samples $\{{\bf x}^{acq}_{i}\}$, we could possibly select very similar views inside the acquisition set $\{{\bf x}^{acq}_{i}\}$ as we do not consider the mutual information between our selections. However, we would face a combinatorial explosion if we directly attempt to maximize the expected information gain between training and acquisition samples and simultaneously minimize the mutual information across acquisition samples. Therefore, we employ a greedy optimization algorithm as illustrated in Alg. 1, which is $\nicefrac{{1}}{{e}}$-approximate in Fisher Information. When batch size $B$ is 1, our algorithm is equivalent to Eq. (8). Please note that we focus on batch active view selection instead of dataset subsampling as it is more related to real-world scenarios where we wish to plan a trajectory for an agent to acquire more training views.

### v2

# FisherRF: Active View Selection and Mapping with Radiance Fields using Fisher Information

@@ Abstract
This study addresses the challenging problem of active view selection and uncertainty quantification within the domain of Radiance Fields. Neural Radiance Fields (NeRF) have greatly advanced image rendering and reconstruction, but the cost of acquiring images poses the need to select the most informative viewpoints efficiently. Existing approaches depend on modifying the model architecture or hypothetical perturbation field to indirectly approximate the model uncertainty. However, selecting views from indirect approximation does not guarantee optimal information gain for the model. By leveraging Fisher Information, we directly quantify observed information on the parameters of Radiance Fields and select candidate views by maximizing the Expected Information Gain (EIG). Our method achieves state-of-the-art results on multiple tasks, including view selection, active mapping, and uncertainty quantification, demonstrating its potential to advance the field of Radiance Fields.
@@ 1 Introduction
Figure 1: A brief overview of our method Given a Radiance Field that was trained with a limited number of views, our method could find the next best view that could maximize information gain by computing the Fisher Information of the radiance field. We illustrate the Information Gain as a heat map on the viewing sphere and show four of the candidate views. Our model can quantify pixel-wise uncertainty, visualized at the bottom left, by examining the Fisher Information on the related parameters of the ray. Our algorithm can be also adapted into an active mapping system, as showcased at the right, that could actively explore and reconstruct the environment.
@@
Neural Radiance Fields brought back image rendering and reconstruction from multiple views to the center of attention in the field of computer vision. Novel volumetric representations of radiance fields and differentiable volumetric rendering enabled unprecedented advances in image-based rendering of complex scenes both in terms of perceptual quality and speed. Neural Radiance Fields (NeRF) have recently garnered significant attention in the field of Computer Vision and have seen enhancements for faster rendering and training, larger-scale reconstruction, dynamic modeling, and generative capabilities.
@@
Recently, 3D Gaussian Splatting [25] has demonstrated distinct advantages in real-time rendering and explicit point-based parameterizations without neural representations. However, to achieve satisfactory rendering quality, numerous viewpoints are required to train a radiance field, especially in large-scale scenarios. It is crucial to establish a criterion for selecting information-maximizing views before obtaining image observations at those locations. Quantifying the observed information of a Radiance Field model is challenging, given that Radiance Field models are typically regression-based and scene-specific. The challenge intensifies when we aim to leverage quantified observed information for active view selection and mapping, especially when the selection candidates are only $SE(3)$ camera poses for acquiring new observations, a.k.a capturing new images.
@@
Previous approaches select viewpoints by quantifying the uncertainty in Radiance Fields. They can be broadly categorized into two groups: variational white-box models and black-box models. White-box models integrate conventional NeRF architectures with Bayesian models, such as reparameterization [43, 36, 48] and Normalizing Flows [55]. Black-box methods, on the other hand, do not modify the existing model architecture but seek to quantify predictive uncertainty by examining the distribution of predicted outcomes [71, 31, 67, 43]. White-box models depend on specific model architectures and are often characterized by slower training times due to the challenges associated with probabilistic learning. Conversely, existing black-box models either focus solely on studying the uncertainty on NeRF-style query points through network prediction [43] or hypothetical perturbation field [16], or rely on Monte Carlo sampling techniques [67] to quantify the uncertainty at the candidate views.
In this study, our primary objective is to quantify the observed information of a Radiance Field model and utilize it to select the optimal view with the highest information gain for downstream tasks, Fig. 1. To achieve this, we propose using Fisher Information, which represents the expectation of observation information. This quantity is directly linked to the second-order derivatives or Hessian matrix of the loss function involved in optimizing Radiance Field models. Importantly, the Hessian of the objective function in volumetric rendering is independent of ground truth data or the actual image measurement. This property allows us to compute the information gain between the training dataset and the candidate viewpoint pool using only the camera parameters of the candidate views. This capability facilitates an efficient next-best-view selection algorithm. In addition to the formulation of Information Gain, the Hessian of the loss function has an intuitive interpretation: the perturbation at flat minima of the loss function. From an intuitive standpoint, one can perceive the Fisher Information matrix as a metric of the curvature of the log-likelihood function at specific parameter instantiations denoted as ${\bf w}^{*}$. Lower Fisher Information suggests that the log-likelihood function exhibits a flatter profile around ${\bf w}^{*}$, implying that the loss is less prone to changes when ${\bf w}^{*}$ is perturbed. The flat minimum interpretation has attracted substantial attention and research within various domains of machine learning [19, 28, 20, 58, 34]. Moreover, as we have estimated the Fisher information for the model’s parameters, which correspond to specific 3D locations in recent Radiance Field models such as 3D Gaussian Splatting [25] and Plenoxels [52], we can derive pixel-wise uncertainty in the model’s predictions by examing the Fisher Information on parameters that contribute to the prediction for each pixel.
@@
We implement the computation of Fisher Information on top of two types of Radiance Field models: point-based 3D Gaussian Splatting [25] and Plenoxels [52]. In 3D Gaussian Splatting, we compute the Hessians of the parameters on each 3D Gaussian with respect to the negative log-likelihood. As Fisher Information is additive, we further extend our view selection algorithm into batch selection and path selection, which are closer to real-world applications. To make the computation of information gain tractable, we use an approximation of the decrease in entropy by its upper bound, which is the trace of the product of the Hessian of the candidate view times the Hessian with respect to all previous views. These matrices are highly sparse due to the disentangling of scene parameters with respect to disparate rays. The sparsity allows us a computation of the matrix trace above that is as cost-effective as back-propagation, enabling us to evaluate views at 70 fps when we use 3D Gaussian Splatting. We carried out extensive evaluations in different benchmarks, including active view selection, active mapping, and uncertainty quantification. The quantitative and qualitative results unequivocally demonstrate that our approach surpasses previous methods and heuristic baseline by a significant margin. To summarize, our main contributions are as follows:
  • A novel formulation to quantify the observed information and Information Gain for Radiance Fields.
@@
  • An effective and efficient view selection method exploiting the sparse structure of the scene rendering problem.
@@
  • An efficient pixel-wise uncertainty quantification and visualization method for 3D Gaussian Splatting.
@@
  • Extensive comparative study showing that our method outperforms existing active approaches on view selection, active mapping, and uncertainty quantification.
@@ 2 Related Works
Active Learning and Radiance Fields are prosperous research fields with various research directions. In this section, we limit our literature review to works that directly relate to our tasks. We refer the readers to literature reviews if they are interested in Radiance Fields [13, 10] or Deep Active Learning [50, 72]. Besides, Fisher Information has been extensively studied in deep active learning [1, 2, 29, 27]. Notably, Kirsch et al. [28] unified existing works in active learning for deep learning problems from the perspective of Fisher Information, which shares many insights with our method.
@@
In the Neural Radiance Fields, Lee et al. [31], Yan et al. [66] and Zhan et al. [71] attempted to quantify the uncertainty in a scene by the distribution of densities on a casted ray. Shen et al. [55, 56] designed Bayesian models by re-parametrizing the NeRF model. However, they only tackled the predictive uncertainty of the model that did not relate to the observed information of the parameters. Sunderhauf et al. [61] propose an additional uncertainty measure in uncharted regions, determined by the ray termination probability on the learned geometry. Regarding concurrent work, Goli et al. [16] introduces a hypothetical perturbation field and interpolates uncertainty over NeRF-style query points for uncertainty quantification and NeRF clean-up task. The perturbation field with respect to query points is also not compatible with recent radiance field models like 3D Gaussian Splatting [25], which do not take query points as inputs. Besides, their algorithm is not designed for an active learning problem since selecting candidates with the largest per-pixel uncertainty may not be optimal, as the information gain for the model parameters is not concerned. In contrast, we focus on active learning problems. We derive our theory from the expected information gain for the model and compute the Hessian for the model parameters, allowing us to apply our algorithm to recent advancements in radiance fields such as 3D Gaussian splatting. Our paper focuses on active view selection and mapping based on our novel expected information gain objective. Moreover, their implementation relies on multiple function calls to PyTorch backward engine whereas our efficient CUDA implementation for the diagonal Hessian computation consumes much less time (11.3 ms $\pm$ 33.3 $\mu$s v.s. 1.1 s $\pm$ 2.5 ms). To the best of our knowledge, we are the first method that quantifies the uncertainty directly on model parameters for Radience Fields, thanks to the powerful framework of 3D Gaussian and our efficient formulation.
Active View Selection and Mapping
Although the next best view selection problem was extensively studied before radiance fields gained popularity, there has not been much literature on active “training” view selection for volumetric rendering. Jin et al. [22] trains a neural network to predict per-pixel uncertainty for view selection. However, their method requires an image capture as the input of the network, which is closer to dataset subsampling rather than active learning. ActiveNeRF [43] studied the next best view selection by directly adding variances as the output of the vanilla NeRF, and it is the closest approach to the view selection problem we are working on. In a larger scope, Active Perception [3, 4] has been studied along with the advancement of robotics and computer vision. Traditionally, frontier-based approach [64] and Rapidly exploring Random Tree (RRT) [30] are the most representative and have been extended and enhanced for different data representations [74, 23, 12, 57] and complex conditions [68, 44]. For the larger scope of Active Simultaneous Localization and Mapping (SLAM), please refer to literature reviews [33, 45]. In recent years, Guedon et al. [17, 18] attempted to improve the coverage for mapping systems using point cloud representation with neural networks. Dhami et al. [11] studied the next best view selection in view planning by first completing the point cloud from partial observation. Georgakis et al. [14] measured uncertainty through deep ensemble. Chaplot et al. [8] and Ramakrishnan et al. [46] developed active exploration methods with deep reinforcement learning. The mapping system that uses implicit feild [75, 59, 42, 51] espeically 3D Gaussians [24, 65, 37, 70] as representation has been widely studied. This presents the need for active mapping algorithms that could maximize the mapping efficiency in an embodied robot system. Similar to ActiveNeRF [43], Ran et al. [48] performed trajectory planning for object-centric autonomous reconstruction systems by predicting variance as an extra output of the network. Yan et al. [67] discussed the intuition of flat minimum and quantified the predictive uncertainty of neural mapping models through the lens of loss landscape. However, they only approximate the uncertainty of the variances from the output of neighboring points. Our method directly quantifies the observed information of the parameters by computing the Hessian matrix of the log-likelihood function of the model’s objective. Besides, we introduce the Expected Information Gain for mapping systems from the first principle as a complementary method to previous heuristics.
@@ 3 Technical Approach
In this section, we first introduce how we use Fisher Information in Sec. 3.1. By leveraging Fisher Information, we demonstrate how to select the next best view and next best batch of views in Sec. 3.2 and Sec. 3.3. We then extend our method to large-scale scenes with active mapping. Finally, we showcase that our method could quantify the pixel-wise uncertainties in Sec. 3.5.
@@ 3.1 Fisher Information in Volumetric Rendering
Fisher Information is a measurement of information that an observation $({\bf x},{\bf y})$ carries about the unknown parameters ${\bf w}$ that model $p({\bf y}|{\bf x};{\bf w})$. In the problem of novel view synthesis, $({\bf x},{\bf y})$ are the camera pose ${\bf x}$ and image observation ${\bf y}$ at pose ${\bf x}$, respectively, whereas ${\bf w}$ are the parameters of the radiance field. The objective of neural rendering is equivalent to minimizing the negative log likelihood (NLL) between rendered images and ground truth images on the holdout set, which inherently represents the quality of scene reconstruction
@@
$\mathcal{I}({\bf w})=-\mathbb{E}_{\log p({\bf y}|{\bf x},{\bf w})}\big{[}\frac{\partial^{2}\log p({\bf y}|{\bf x},{\bf w})}{\partial{\bf w}^{2}}\big{|}{\bf w}\big{]}={\mathbf{H}}^{\prime\prime}[{\bf y}|{\bf x},{\bf w}]$ (2)
@@ 3.2 Next Best View Selection Using Fisher Information
$\mathcal{I}[{\bf w}^{*};\{{\bf y}^{acq}_{i}\}|\{{\bf x}^{acq}_{i}\}]$
$=\mathbb{H}[{\bf w}^{*}|D^{train}]-\mathbb{H}[{\bf w}^{*}|\{{\bf y}^{acq}_{i}\},\{{\bf x}^{acq}_{i}\},D^{train}]$ (3)
@@
When the log-likelihood has the form of Eq. (1), in our case, the rendering error, the difference of the entropies in the R.H.S. of Eq. (3) can be approximated as [28]:
@@ 3.3 Batch Active View Selection
Selecting multiple views to capture new images is useful for its possible applications, such as view planning and scene reconstruction. If we simply use Eq. (5) to select a batch of acquisition samples $\{{\bf x}^{acq}_{i}\}$, we could possibly select very similar views inside the acquisition set $\{{\bf x}^{acq}_{i}\}$ as we do not consider the mutual information between our selections. However, we would face a combinatorial explosion if we directly attempt to maximize the expected information gain between training and acquisition samples and simultaneously minimize the mutual information across acquisition samples. Therefore, we employ a greedy optimization algorithm as illustrated in Alg. 1, which is $\nicefrac{{1}}{{e}}$-approximate in Fisher Information [41]. When batch size $B$ is 1, our algorithm is equivalent to Eq. (5). Please note that we focus on batch active view selection instead of dataset subsampling as it is more related to real-world scenarios where we wish to plan a trajectory for an agent to acquire more training views.